# Extensive entanglement between coupled Tomonaga-Luttinger liquids in and out of equilibrium

### v1

@@ 1 Introduction
One of the main bottlenecks is that most results in the literature of entanglement for 1D systems focus on partitioning across the 1D axis (also known as the transversal cut, see Fig. 1[26, 27, 28, 29, 30, 31], dividing the total length of the system into non-overlapping segments. It is a ubiquitous expectation that the ground state entanglement entropy for this partition follows the area law in the presence of a spectral gap [30], which in 1D implies that the entanglement does not grow with longer subsystem length; for a gapless system the entanglement grows with subsystem’s length albeit only logarithmically [29]. While parallel 1D Bose gases experiments confirm the area law for mutual information at finite temperatures [19], entanglement with transversal partition has not been reported, possibly because only microscopic amounts can be naturally generated and it is quickly diluted by thermal noise [19, 5, 6, 7, 8].
@@
In this work, we highlight a better strategy to observe entanglement, i.e. across the longitudinal cut (dashed line in Fig. 1) which partitions the system along the 1D axis, dividing it into two transversally separated 1D gases. This partitioning is the more natural strategy for experimentally observing entanglement, since the mechanisms for generating strong correlations are well understood. For example, correlation can be introduced via tunnel-coupling interaction, implemented by adjusting the height of the double-well barrier (Fig. 1a) [19, 11, 18, 12], or by coherently splitting the potential from a single to double-well (Fig. 1b) [17, 16, 32, 33, 34, 35]. Indeed, recent experiments have witnessed entanglement in the longitudinal partition through number squeezing [36]. Previously, progress was limited by a lack of experimental readout techniques, but recent advances including outcoupling measurement [37] and common phase interferometry [38] have overcome this barrier, providing new experimental windows to probe entanglement. However, surprisingly little is known about the theory behind entanglement in the longitudinal partition.
@@
Beyond its experimental relevance, analyzing the longitudinal cut is also of theoretical interest. It enables quantum simulation of entanglement between multiple quantum fields coexisting and interacting in a common (1+1D) spacetime [39, 40, 41, 42, 43], which is not as well understood as the more conventional spatial partition of a single quantum field. In condensed matter physics, this setting is also relevant for studying systems with ladder geometries [44, 45, 46], particularly to study the interplay between inter-chain and intra-chain coupling. Previous works [47, 48] have shown that the ground state entanglement entropy between two coupled TLLs scales linearly with subsystem’s length, which they characterize as ‘boundary law’. We remark that in this case, boundary law is not synonymous with the standard area law since, due to embedding in higher dimension, the boundary size scales as the subsystem’s volume. Thus, coupled TLLs contain an extensive yet spatially structured entanglement arising from local interactions, which distinguish them from generic many-body excited states also obeying volume law scaling but with a high degree of scrambling [49, 50].
@@
The paper is organized as follows. After this introduction, we present preliminaries in Sec. 2 which include the Hamiltonian of coupled 1D Bose gases (Subsec. 2.1), covariance matrix for Gaussian fields (Subsec. 2.2), and procedure to compute entanglement and mutual information of mixed Gaussian states (Subsec. 2.3). In Sec. 3, we present our results on logarithmic negativity and mutual information of tunnel-coupled system at finite temperatures. Then, in Sec. 4 we extend our analysis to a non-equilibrium state by considering the system after coherent splitting. Finally, we present the summary and outlook in Sec. 5.
@@ 2.1 Coupled 1D Bose gases Hamiltonian
The low-energy Hamiltonian governing the dynamics of these fluctuations is given by the Tomonaga-Luttinger liquid (TLL) model, which in the quasicondensate regime is given by [52]
@@
We are interested in the entanglement between the two TLLs. A common way to introduce correlation between them is to add a tunnel-coupling of the form $\hbar J(\hat{\psi}^{A\dagger}\hat{\psi}^{B}+\text{h.c.})$ to the Hamiltonian, with $J$ being the tunneling strength. This interaction is experimentally realized by adjusting the height of the double-well barrier. Such tunneling induces a non-linear interaction of the form
@@
in the low-energy model, assuming small density fluctuations $\delta\hat{n}^{a}(z)\ll n_{\rm 1D}$.
@@
In Eq. (6), we used a harmonic approximation for the cosine term valid in the strong coupling regime (when $J$ dominates over other couplings), giving a Gaussian model for the coupled systems. In this paper, we focus on this Gaussian limit where tunneling only adds a mass term to the antisymmetric sector, realizing the Klein-Gordon field [55, 19, 24].
The underlying tunnel-coupled Hamiltonian can be diagonalized using a Bogoliubov transformation, resulting in two-species bosonic theory comprising a massless mode (++) and a massive (-) mode. The diagonalized Hamiltonian of the non-zero momentum modes is given by
@@
If the system is partitioned into the symmetric and antisymmetric fields, as usually done in the literature, there would be no correlation between the two sectors in thermal equilibrium. But our interest lies in the longitudinal cut shown in Fig. 1, where the system is partitioned into gas $A$ and gas $B$. With this choice, the fields remain correlated at thermal equilibrium as long as $J>0$. Even without any tunneling ($J=0$), correlation can be generated in the initial state by pushing the system far away from equilibrium, e.g., by coherently splitting the system from a single to double-well [14].
@@ 2.2 Covariance matrix for Gaussian fields
where the allowed momenta are $k=N\pi/L$ with $N=1,2,3,\ldots,\Lambda$ and the positive integer $\Lambda$ is chosen such that $k_{\Lambda}=\Lambda\pi/L\sim\xi_{h}^{-1}$. The prefactor $\sqrt{2/L}$ ensures proper commutation relation $\delta\hat{n}^{a}_{k},\hat{\phi}^{b}_{q}]=i\delta_{kq}\delta_{ab}$. From here onwards, we consider the case where the zero-mode $\hat{\phi}_{0}^{a},\delta\hat{n}_{0}^{a}$ are removed from the data by phase referencing and adjusting for atom number fluctuations.
The covariance matrix $\Gamma\in\mathbb{R}^{4\Lambda\times 4\Lambda}$ is then constructed from the two-point correlation functions of the quadrature operator $\hat{Q}^{a}$ defined as
@@ 2.3 Entanglement measure and mutual information for Gaussian states
$E_{\mathcal{N}}\equiv\sum_{q=1}^{2\Lambda}\max\left\{0,-\log\left(2\gamma_{q}^{\top_{B}}\right)\right\}>0$ (16)
@@ 3 Results: Entanglement and mutual information at finite temperatures
In tunnel-coupled Luttinger liquids, the low-energy Hamiltonian separates into symmetric (++) and antisymmetric sectors (-). The system is thus described by a two-species bosonic theory comprising a massless symmetric mode and a massive antisymmetric mode. Assuming Neumann boundary conditions, we use Bogoliubov transformation [56, 57, 51, 18] to expand the phase and density fluctuations in terms of bosonic creation and annihilation operators $\hat{b}_{k}^{\pm}$ in the two sectors
@@
where $E_{k}=(\hbar k)^{2}/(2m)$ is the free particle dispersion, $\varepsilon_{k}^{\pm}$ are the Bogoliubov energy spectra defined in Eq. (8).
In this section, we will assume the system is at thermal equilibrium, having a density matrix $\rho\propto\exp(-\beta\hat{H)}$ with inverse temperature $\beta$. The elements of the covariance matrix $\Gamma^{ab}$ can be calculated from Eqs. (18) and (19). For example, the phase-phase correlations are
@@
$E_{\mathcal{N}}=\sum_{k>0}^{k_{\Lambda}}\max\left\{0,-\log\left[\sqrt{\frac{\varepsilon_{k}^{-}}{\varepsilon_{k}^{+}}\frac{E_{k}}{E_{k}+2\hbar J}(1+2\eta_{k}^{+})(1+2\eta_{k}^{-})}\right]\right\},$ (21)
where
$\eta_{k}^{\pm}=\braket{(\hat{b}_{k}^{\pm})^{\dagger}\hat{b_{k}^{\pm}}}=[\exp(\beta\varepsilon_{k}^{\pm})-1]^{-1}$ (22)
are the mean thermal occupations in the symmetric and antisymmetric sectors. The result can be generalized to the case where each sector has a different temperature $(\beta^{+}\neq\beta^{-})$ by appropriately defining $\eta_{k}^{\pm}$. From Eq. (21), we see a competition between tunneling strength $J$ and the temperature $T=(k_{B}\beta)^{-1}$ in establishing entanglement ($k_{B}$ is the Boltzmann constant). On the one hand, tunneling tends to create entanglement since the spectral ratio satisfies $(\varepsilon_{k}^{-}/\varepsilon_{k}^{+})(E_{k}/(E_{k}+2\hbar J))\leq 1$ with the inequality being saturated at $J=0$. On the other hand, temperature tends to destroy entanglement with the term $(1+2\eta_{k}^{+})(1+2\eta_{k}^{-})\geq 1$.
@@
$T^{*}=\sup_{0<k\leq k_{\Lambda}}\left\{T_{k}^{*}\Big{|}\tanh\left(\frac{\varepsilon_{k}^{+}}{2k_{B}T^{*}_{k}}\right)\tanh\left(\frac{\varepsilon_{k}^{-}}{2k_{B}T_{k}^{*}}\right)=\frac{\varepsilon_{k}^{-}}{\varepsilon_{k}^{+}}\frac{E_{k}}{E_{k}+2\hbar J}\right\},$ (23)
easily solvable by root-finding. Applying this formula reveals that the threshold temperature increases with stronger tunneling (Fig. 2a). Similarly, one may fix the temperature and find the critical tunneling strength needed to observe entanglement, which would be stronger for higher temperatures.
@@
At this point, we remark that while the Hamiltonian (6) can be written as coupled quantum harmonic oscillators, whose thermal entanglement and threshold temperature have been studied in a general setting [8, 7], we identify a solvable and experimentally relevant instance of such models. We also go further by evaluating the experimental attainability of the threshold temperature and exploring the entanglement scaling with tunable system parameters such as tunneling strength, system’s length, and mean density. Later in the section, we will probe the transition point between quantum and classical correlation by comparing logarithmic negativity with mutual information.
@@
Figure 2: Entanglement at finite temperatures for tunnel-coupled gases.— (a) Entanglement phase diagram with respect to tunneling strength $\tilde{J}=J/(2\pi)$ and temperature $T$ as measured by logarithmic negativity $E_{\mathcal{N}}$. The dashed line is the threshold temperature $T^{*}$ above which entanglement vanishes ($E_{\mathcal{N}}=0$) for all modes $k\leq k_{\Lambda}$. The dashed-dotted line is the first threshold temperature $T^{*}_{k_{1}}$ above which the entanglement in the first mode $k_{1}=\pi/L$ vanishes, while still possibly present in higher momentum modes with $k>k_{1}$. (b) The decay of $E_{\mathcal{N}}$ with increasing $T$ for fixed tunneling $\tilde{J}=5\;\rm Hz$. For panels (b)-(d), the solid lines are calculated with our analytical formula (21), while the triangles are numerically computed with numerical diagonalization. The dashed and dashed-dotted lines have the same meaning as in (a). (c) Linear scaling of $E_{\mathcal{N}}\propto L$ for fixed temperature $T=2\;\rm nK$ and fixed tunneling strength $\tilde{J}=5\;\rm Hz$. The cutoff $\Lambda$ changes for different $L$ since we set $k_{\Lambda}=\Lambda\pi/L\sim\xi_{h}^{-1}$. In panels (a) - (c), mean density is fixed at $n_{\rm 1D}=75\;\rm\mu m^{-1}$, but the results do not significantly change for other $n_{\rm 1D}$ values. (d) Sub-linear scaling of $E_{\mathcal{N}}$ with mean density $n_{\rm 1D}$. The difference between scaling behavior in (c) and (d) underscores the collective nature of 1D excitations. For panels (a), (b), and (d), the condensate length is fixed at $L=100\;\rm\mu m$. The mass $m$ is the atomic mass 87Rb. The interaction strength $g$ is close to the value reported in the experiments (see e.g. Refs. [19, 25] among others) and fixed throughout the paper, i.e. $g\approx 2\hbar a_{s}\omega_{\perp}$ where $a_{s}\approx 5.2\;\rm nm$ is the scattering length, and $\omega_{\perp}=2\pi\times 2\;\rm kHz$ is the transverse trapping frequency.
@@
We plot a phase diagram for entanglement as measured by $E_{\mathcal{N}}$ with varying temperature $T$ and tunneling strength $J$ in Fig. 2a. We identify the parameter regimes where the system is entangled ($E_{\mathcal{N}}>0$) and where it is separable ($E_{\mathcal{N}}=0$), separated by the threshold temperature $T^{*}$ curve. In the entangled regime, entanglement intuitively increases with stronger tunneling and decreases with increasing temperature. For the experimental parameters used in Refs.   [18, 19, 25], $T^{*}$ lies in the range of $1\sim 5\;\rm nK$, about one order of magnitude lower than the temperatures achieved in these experiments.
@@
Given a fixed tunneling strength, the largest $T_{k}^{*}$ typically occurs for the largest accessible mode $k_{\Lambda}$. We interpret this as the entanglement in higher momentum modes being more robust to thermal noise. As one raises the temperature, thermal noise corrupts the entanglement in all modes, but it corrupts the entanglement in low momentum modes more quickly. For instance, above the first threshold $T>T^{*}_{k_{1}}$ ($k_{1}=\pi/L$), the entanglement between $A$ and $B$ vanishes for $k_{1}$, while still possibly present for $k>k_{1}$ (Fig. 2b). This process continues until all entanglement in the modes $k\leq k_{\Lambda}$ vanishes at $T=T^{*}$ and we no longer observe entanglement. Naturally, the ability to detect entanglement depends on the cutoff $\Lambda$. For a fixed cutoff $k_{\Lambda}\sim\xi_{h}^{-1}$ the system can be regarded as ‘effectively separable’ at $T>T^{*}$, even though entanglement may still persist above the cutoff. Conversely, the experimental imaging resolution is typically lower than $\xi_{h}^{-1}$, and so depending on the specific resolution, the threshold temperature in experiments can be lower than in Fig. 2.
@@
We next investigate the scaling of entanglement with the subsystem’s length $L$. Previous theoretical work [47, 48] has shown that in the ground state, entanglement behaves extensively, meaning it grows linearly with $L$. We obtain the same result for finite temperature states as long as the temperature is sufficiently far away from $T^{*}$ as shown in Fig. 2c. We next provide a simple argument to explain this. Given a fixed resolution, the number of allowed $k$-modes increases linearly with $L$, i.e. in the continuum limit $\sum_{k}\rightarrow(L/2\pi)\int dk$. In our setting, this is the dominant mechanism by which entanglement grows with length $E_{\mathcal{N}}\propto L$. We emphasize that this extensivity arises due to the embedding of the 1D systems in 2D space, so it does not violate the standard area law for a strictly 1D system. Interestingly, the scaling behavior is sublinear if we increase mean density $n_{\rm 1D}$ instead of $L$ (Fig. 2d), underscoring the collective nature of excitations, as entanglement does not just scale linearly with atom number $n_{\rm 1D}L$.
@@
$I(A:B)=$ $\sum_{k>0}^{k_{\Lambda}}2\left[\left(\lambda_{k}+\frac{1}{2}\right)\log\left(\lambda_{k}+\frac{1}{2}\right)-\left(\lambda_{k}-\frac{1}{2}\right)\log\left(\lambda_{k}-\frac{1}{2}\right)\right]$
@@
where $\lambda_{k}$ are the symplectic eigenvalues of the reduced covariance matrix, which are given by
$\lambda_{k}=\frac{1}{4}\sqrt{\left[(1+2\eta_{k}^{+})+C_{k}(1+2\eta_{k}^{-})\right]\left[(1+2\eta_{k}^{+})+C_{k}^{-1}(1+2\eta_{k}^{-})\right]}$ (25)
and $C_{k}=(\varepsilon_{k}^{-}/\varepsilon_{k}^{+})(E_{k}/(E_{k}+2\hbar J))$ is the spectral ratio also appears in the expression for $E_{\mathcal{N}}$ in Eq. (21).

### v2

@@ 1 Introduction
One of the main bottlenecks is that most results in the literature of entanglement for 1D systems focus on partitioning across the 1D axis (also known as the transversal cut, see Fig. 1) [26, 27, 28, 29, 30, 31], dividing the total length of the system into non-overlapping segments. It is a ubiquitous expectation that the ground state entanglement entropy for this partition follows the area law in the presence of a spectral gap [30], which in 1D implies that the entanglement does not grow with longer subsystem length. For a gapless system the entanglement grows with subsystem’s length albeit only logarithmically [29]. While parallel 1D Bose gases experiments confirm the area law for mutual information at finite temperatures [19], entanglement with transversal partition has not been reported, possibly because only microscopic amounts can be naturally generated and it is quickly diluted by thermal noise [19, 5, 6, 7, 8].
@@
In this work, we highlight a better strategy to observe entanglement, i.e. across the longitudinal cut (dashed line in Fig. 1), which partitions the system along the 1D axis and divides it into two transversely separated 1D gases. This partitioning is the more natural strategy for experimentally observing entanglement, since the mechanisms for generating strong correlations are well understood. For example, correlation can be introduced via tunnel-coupling interaction, implemented by adjusting the height of the double-well barrier (Fig. 1a) [19, 11, 18, 12], or by coherently splitting the potential from a single to double-well (Fig. 1b) [17, 16, 32, 33, 34, 35]. Entanglement between spatially separated (3D) Bose–Einstein condensates (BEC) has been demonstrated [36, 37, 38] using a witness based on the total spin of each atomic cloud. In parallel 1D quasicondensates, similar methods have observed quantum-limited number squeezing [39, 40, 41] which witness many-particle entanglement [42, 43]. In contrast to BECs, finite-momentum excitations play a central role in 1D quasicondensates. A key challenge is to certify and measure entanglement between such finite-momentum excitations across different spatially separated modes [44]. In 1D quasicondensates, the measurements of these finite-momentum excitations are accessible through spatially resolved matter-wave interferometry [45, 46, 47] which enables local measurements of the quantum fields and provides opportunity for probing entanglement in the longitudinal partition.
@@
Beyond its experimental relevance, analyzing the longitudinal cut is also of theoretical interest. It enables quantum simulation of ‘field space entanglement’ [48, 49, 50, 51, 52], i.e. entanglement between multiple quantum fields coexisting and interacting in a common (1+1D) spacetime, which is not as well understood as the more conventional spatial partition of a single quantum field. In condensed matter physics, this setting is also relevant for studying ‘bulk entanglement’ in systems with ladder geometries [53, 54, 55, 56]. Previous theoretical works [57, 58] in this direction have shown that the ground state entanglement entropy between two coupled TLLs scales linearly with subsystem’s length, which they characterize as ‘boundary law’. We remark that in this case, boundary law is not synonymous with the standard area law since the boundary size scales as the subsystem’s volume. Thus, coupled TLLs contain an extensive yet spatially structured entanglement arising from local interactions, which distinguish them from generic many-body excited states also obeying volume law scaling but with a high degree of scrambling [59, 60].
@@
The paper is organized as follows. After this introduction, we present preliminaries in Sec. 2 which include the Hamiltonian of coupled 1D Bose gases (Subsec. 2.1), covariance matrix for Gaussian fields (Subsec. 2.2), and procedures to compute entanglement and mutual information of mixed Gaussian states (Subsec. 2.3). In Sec. 3, we present our results on logarithmic negativity and mutual information of tunnel-coupled system at finite temperatures. Then, in Sec. 4 we extend our analysis to a non-equilibrium state by considering the system after coherent splitting. Finally, we present the summary and outlook in Sec. 5.
@@ 2.1 Coupled 1D Bose gases Hamiltonian
The microscopic Hamiltonian of the system is given by the Lieb-Liniger model [62]. At low-energies, it is well described by the Tomonaga-Luttinger liquid (TLL) Hamiltonian of the form [63]
@@
Here, we have assumed explicit forms of the speed of sound and Luttinger parameters valid in the quasicondensate regime. We are interested in the entanglement between the two TLLs. A common way to introduce correlation between them is to add a tunnel-coupling of the form $\hbar J(\hat{\psi}^{A\dagger}\hat{\psi}^{B}+\text{h.c.})$ to the Hamiltonian with $J$ being the tunneling strength. Such tunneling induces a non-linear interaction of the form
@@
in the low-energy model, assuming small density fluctuations $\delta\hat{n}^{a}(z)\ll n_{\rm 1D}$. This interaction can be experimentally realized by adjusting the height of the double-well barrier separating the two quasicondensates.
@@
In Eq. (6), we have used a harmonic approximation for the cosine term valid in the strong coupling regime (when $J$ dominates over other couplings), giving a Gaussian model for the coupled systems. In this paper, we focus on this Gaussian limit where tunneling only adds a mass term to the antisymmetric sector, realizing the Klein-Gordon field [64, 19, 24].
For weak atomic repulsion and at sufficiently low energies, the underlying Hamiltonian can be diagonalized using a Bogoliubov transformation, resulting in two-species bosonic theory comprising a massless mode (++) and a massive (-) mode. The diagonalized Hamiltonian of the non-zero momentum modes is given by
@@
Note that Eq. (8) takes into account deviation from linear dispersion in the phononic regime. If the system is partitioned into the symmetric (++) and antisymmetric (-) fields, as usually done in the literature, there would be no correlation between the two sectors in thermal equilibrium. But our interest lies in the longitudinal cut shown in Fig. 1, where the system is partitioned into gas $A$ and gas $B$. With this choice, the fields remain correlated at thermal equilibrium as long as $J>0$. Even without any tunneling ($J=0$), correlation can be generated in the initial state by pushing the system far away from equilibrium, e.g., by coherently splitting the trap from a single to double-well [14].
@@ 2.2 Covariance matrix for Gaussian fields
where the allowed momenta are $k=N\pi/L$ with $N=1,2,3,\ldots,\Lambda$ and the positive integer $\Lambda$ is chosen such that $k_{\Lambda}=\Lambda\pi/L\sim\xi_{h}^{-1}$. The prefactor $\sqrt{2/L}$ ensures proper commutation relation $\delta\hat{n}^{a}_{k},\hat{\phi}^{b}_{q}]=i\delta_{kq}\delta_{ab}$.
From here onwards, we focus on correlation in the non-zero momentum modes, e.g. we consider the case where the zero-modes $\hat{\phi}_{0}^{a},\delta\hat{n}_{0}^{a}$ are removed from the data by phase referencing and adjusting for atom number fluctuations. This assumption differentiates our approach from previous studies of particle-entanglement in 1D quasicondensates [40, 68], which relies on probing entanglement through zero-mode squeezing in the relative sector. The relevant covariance matrix $\Gamma\in\mathbb{R}^{4\Lambda\times 4\Lambda}$ is then constructed from the two-point correlation functions of the quadrature operator $\hat{Q}^{a}$ defined as
@@ 2.3 Entanglement measure and mutual information for Gaussian states
$E_{\mathcal{N}}\coloneqq\sum_{q=1}^{2\Lambda}\max\left\{0,-\log\left(2\gamma_{q}^{\top_{B}}\right)\right\}>0$ (16)
@@ 3 Results: Entanglement and mutual information at finite temperatures
Since the low-energy Hamiltonian separates into symmetric (++) and antisymmetric sectors (-), the system is described by a two-species bosonic theory comprising a massless symmetric mode and a massive antisymmetric mode. Assuming Neumann boundary conditions, we use Bogoliubov transformation [65, 66, 61, 18] to expand the phase and density fluctuations in terms of bosonic creation and annihilation operators $\hat{b}_{k}^{\pm}$ in the two sectors
@@
where $E_{k}=(\hbar k)^{2}/(2m)$ is the free particle dispersion, $\varepsilon_{k}^{\pm}$ are the Bogoliubov energy spectra defined in Eq. (8). In this section, we will assume the system is at thermal equilibrium, having a density matrix $\rho\propto\exp(-\beta\hat{H)}$ with inverse temperature $\beta$. The elements of the covariance matrix $\Gamma^{ab}$ can be calculated from Eqs. (18) and (19). For example, the phase-phase correlations are
@@
$E_{\mathcal{N}}=\sum_{k>0}^{k_{\Lambda}}\max\left\{0,-\log\left[\sqrt{C_{k,J}(1+2\eta_{k}^{+})(1+2\eta_{k}^{-})}\right]\right\},$ (21)
where the spectral ratio $C_{k,J}$ is defined as
$C_{k,J}\coloneqq\frac{\varepsilon_{k}^{-}}{\varepsilon_{k}^{+}}\frac{E_{k}}{E_{k}+2\hbar J},$ (22)
and
$\eta_{k}^{\pm}\coloneqq\braket{(\hat{b}_{k}^{\pm})^{\dagger}\hat{b_{k}^{\pm}}}=[\exp(\beta\varepsilon_{k}^{\pm})-1]^{-1}$ (23)
are the mean thermal occupations in the symmetric and antisymmetric sectors. The result can be generalized to the case where each sector has a different temperature $(\beta^{+}\neq\beta^{-})$ by appropriately defining $\eta_{k}^{\pm}$. From Eq. (21), we see a competition between tunneling strength $J$ and the temperature $T=(k_{B}\beta)^{-1}$ in establishing entanglement ($k_{B}$ is the Boltzmann constant). On the one hand, tunneling tends to create entanglement since the spectral ratio satisfies $C_{k,J}\leq 1$ with the inequality being saturated at $J=0$. This implies that at zero temperatures ($\eta_{k}^{+}=\eta_{k}^{-}=0$), all modes are entangled as long as $J>0$. However, as expected, temperature tends to destroy entanglement with the term $(1+2\eta_{k}^{+})(1+2\eta_{k}^{-})\geq 1$.
@@
$T^{*}=\sup_{0<k\leq k_{\Lambda}}\left\{T_{k}^{*}\Big|\tanh\left(\frac{\varepsilon_{k}^{+}}{2k_{B}T^{*}_{k}}\right)\tanh\left(\frac{\varepsilon_{k}^{-}}{2k_{B}T_{k}^{*}}\right)=C_{k,J}\right\},$ (24)
@@
easily solvable by root-finding. At this point, we remark that while the Hamiltonian (6) can be written as coupled quantum harmonic oscillators, whose thermal entanglement and threshold temperature have been studied in a general setting [8, 7], we identify a solvable and experimentally relevant instance of such models. We also go further by evaluating the experimental attainability of the threshold temperature and exploring the entanglement scaling with tunable system parameters such as tunneling strength, system’s length, and mean density. Later in the section, we will probe the transition point between quantum and classical correlation by comparing logarithmic negativity with mutual information.
@@
Figure 2: Entanglement at finite temperatures for tunnel-coupled gases.— (a) Entanglement phase diagram with respect to tunneling strength $\tilde{J}=J/(2\pi)$ and temperature $T$ as measured by logarithmic negativity $E_{\mathcal{N}}$. The dashed line is the threshold temperature $T^{*}$ above which entanglement vanishes ($E_{\mathcal{N}}=0$) for all modes $k\leq k_{\Lambda}$. The dashed-dotted line is the first threshold temperature $T^{*}_{k_{1}}$ above which the entanglement in the first mode $k_{1}=\pi/L$ vanishes, while still possibly present in higher momentum modes with $k>k_{1}$. (b) The decay of $E_{\mathcal{N}}$ with increasing $T$ for fixed tunneling $\tilde{J}=5\;\rm Hz$. For panels (b)-(d), the solid lines are calculated with our analytical formula (21), while the triangles are computed with numerical diagonalization. The dashed and dashed-dotted lines have the same meaning as in (a). (c) Linear scaling of $E_{\mathcal{N}}\propto L$ for fixed temperature $T=2\;\rm nK$ and fixed tunneling strength $\tilde{J}=5\;\rm Hz$. The cutoff $\Lambda$ changes for different $L$ since we set $k_{\Lambda}=\Lambda\pi/L\sim\xi_{h}^{-1}$. In panels (a) - (c), mean density is fixed at $n_{\rm 1D}=75\;\rm\mu m^{-1}$, but the results do not significantly change for other $n_{\rm 1D}$ values. (d) Sub-linear scaling of $E_{\mathcal{N}}$ with mean density $n_{\rm 1D}$. The difference between scaling behavior in (c) and (d) underscores the collective nature of 1D excitations. For panels (a), (b), and (d), the condensate length is fixed at $L=100\;\rm\mu m$. The mass $m$ is the atomic mass 87Rb. The interaction strength $g$ is close to the value reported in the experiments (see e.g. Refs. [19, 25] among others) and fixed throughout the paper, i.e. $g\approx 2\hbar a_{s}\omega_{\perp}$ where $a_{s}\approx 5.2\;\rm nm$ is the scattering length, and $\omega_{\perp}=2\pi\times 2\;\rm kHz$ is the transverse trapping frequency.
@@
We plot a phase diagram for entanglement as measured by $E_{\mathcal{N}}$ with varying temperature $T$ and tunneling strength $J$ in Fig. 2a. We identify the parameter regimes where the system is entangled ($E_{\mathcal{N}}>0$) and where it is separable ($E_{\mathcal{N}}=0$), separated by the threshold temperature $T^{*}$ curve. In the $E_{\mathcal{N}}>0$ regime, entanglement intuitively increases with stronger tunneling and drops with rising temperature. The threshold temperature $T^{*}$ is also found to monotonically increase with stronger tunneling. For the experimental parameters used in Refs. [18, 19, 25], $T^{*}$ lies in the range of $1\sim 5\;\rm nK$, about one order of magnitude lower than the temperatures achieved in these experiments. The results do not significantly change if we allow for different temperatures in symmetric and antisymmetric sectors ($T^{+}\neq T^{-}$), both temperatures need to be of order $1\sim 5\;\rm nK$ to observe entanglement.
@@
For a fixed tunneling strength, the largest $T_{k}^{*}$ typically occurs for the mode with the highest accessible momentum $k_{\Lambda}$. We interpret this as the entanglement in higher momentum modes being more robust to thermal noise. As one raises the temperature, thermal noise corrupts the entanglement in all modes, but it corrupts the entanglement in low momentum modes more quickly. For instance, above the first threshold $T>T^{*}_{k_{1}}$ ($k_{1}=\pi/L$), the entanglement between $A$ and $B$ vanishes for $k_{1}$, while still possibly present for $k>k_{1}$ (Fig. 2b). This process continues until all entanglement in the modes $k\leq k_{\Lambda}$ vanishes at $T=T^{*}$ and we no longer observe entanglement. Naturally, the ability to detect entanglement depends on the cutoff $\Lambda$. For a fixed cutoff $k_{\Lambda}\sim\xi_{h}^{-1}$ the system can be regarded as ‘effectively separable’ at $T>T^{*}$, even though entanglement may still persist above the cutoff. Conversely, the experimental imaging resolution is typically lower than $\xi_{h}^{-1}$, and so depending on the specific resolution, the threshold temperature in experiments can be lower than in Fig. 2.
@@
We next investigate the scaling of entanglement with the subsystem’s length $L$. Previous theoretical works [57, 58] have shown that in the ground state, entanglement behaves extensively, meaning it grows linearly with $L$. We obtain the same result for finite temperature states as long as the temperature is sufficiently far away from $T^{*}$ as shown in Fig. 2c. We next provide a simple argument to explain this. Given a fixed resolution, the number of allowed $k$-modes increases linearly with $L$, i.e. in the continuum limit $\sum_{k}\rightarrow(L/2\pi)\int dk$. In our setting, this is the dominant mechanism by which entanglement grows with length $E_{\mathcal{N}}\propto L$. We emphasize that this extensivity arises due to the embedding of the 1D systems in 2D space, so it does not violate the standard area law for a strictly 1D system. Interestingly, the scaling behavior is sublinear if we increase mean density $n_{\rm 1D}$ instead of $L$ (Fig. 2d), underscoring the collective nature of excitations, as entanglement does not just scale linearly with atom number $n_{\rm 1D}L$. It also highlights the importance of interatomic interaction for observing the entanglement since $n_{\rm 1D}$ appears in Eq. (21) only through the interaction energy $2gn_{\rm 1D}$ in the Bogoliubov spectra $\varepsilon_{k}^{\pm}$ [Eq. (8)].
@@
$I(A:B)=$ $\sum_{k>0}^{k_{\Lambda}}2\left[\left(\lambda_{k,J}+\frac{1}{2}\right)\log\left(\lambda_{k,J}+\frac{1}{2}\right)-\left(\lambda_{k,J}-\frac{1}{2}\right)\log\left(\lambda_{k,J}-\frac{1}{2}\right)\right]$
@@
where $\lambda_{k,J}$ are the symplectic eigenvalues of the reduced covariance matrix given by
$\lambda_{k,J}=\frac{1}{4}\sqrt{\left[(1+2\eta_{k}^{+})+C_{k,J}(1+2\eta_{k}^{-})\right]\left[(1+2\eta_{k}^{+})+C_{k,J}^{-1}(1+2\eta_{k}^{-})\right]},$ (26)
and $C_{k,J}$ is the spectral ratio defined in Eq. (22). Note that the mutual information in Eq. (25) reduces to twice the entanglement entropy in the limit of zero temperature ($\eta_{k}^{+}=\eta_{k}^{-}=0$).